29. What is the product of the complex numbers (-31 + 4) and (31 + 4) ?
Answer #1

(-31+4). (31+4) =(4-3i)(4+3i) = 42-(33)? = 16-912 = 16+9= 25 a product of -3i+4, 3ith = [25